34. Suppose autonomous imports exceeds exports. In the Keynesian model, when we add the
negative net exports to C + I + G …
[1] the A curve shifts upwards
[2] the A curve shifts downwards
[3] the A curve does not shift
[4] the multiplier increases
35. Nthabiseng’s monthly disposable income increases from R3 000 to R3 400. As a result, her
monthly savings increase from R400 to R560. This implies that her marginal propensity to
consume is …
[1] 0,40.
[2] 0,20.
[3] 0,60.
[4] 0,80.
1
Expert's answer
2020-10-05T14:26:05-0400

34. Answer: [1] the A curve shifts upwards

Explanation:

If the imports increase the aggregate demand in the domestic economy would decrease because the amount of money is being spent on imports. The benefit of this demand which is being met with imports would go to the country from where it is being imported.the employment in that country would increase.


35. Answer: [3] 0,60.

Solution:

Her marginal propensity to save=Change in savings/change in income

=160/400

=0.40

Given that: MPC=1-MPS

=1-0.40

=0.60